OpenStudy (anonymous):

Vertify the identity. sin^2 (θ−π/2) = cos^2θ

OpenStudy (anonymous):

\[ \sin^2(\theta - \frac \pi 2)= = \left(- \sin( \theta - \frac \pi 2)\right)^2= \sin^2( \theta - \frac \pi 2) \]

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

Is it (-cosθ) (-cosθ) = cos^2θ

OpenStudy (anonymous):

Yes